OpenStudy (anonymous):

A geometric sequence has first term 6 and common ratio 1.5. The sum of the first n terms of the sequence is 79.125. Find n.(please help me. Willing to give a medal)

OpenStudy (anonymous):

thanks @cj49

OpenStudy (campbell_st):

the information is correct but the format of the fomula is incorrect... which will lead to an incorrect solution you need \[79.125 = \frac{6( 1 - (1.5)^n}{(1 - 1.5)}\] the common ratio needs to be in the same position in the numerator and denominator

OpenStudy (anonymous):

it says that I have to find n. and n= 5. but i dont know how to find that

OpenStudy (campbell_st):

ok... so you will need to use logarithms to find n. an alternative is guess and check...

OpenStudy (campbell_st):

@cj49 it is

OpenStudy (cj49):

n = the number of terms that means there are total 5 terms

OpenStudy (cj49):

6(1.5^n-1) 6(1.5^1)=2nd term 6(1.5^2)=3rd term 6(1.5^3)=4th term 6(1.5^4)=5th term

OpenStudy (campbell_st):

do if you do it correctly you get \[\frac{79.125 \times (1 - 1.5)}{6} =-7.59325\] which means \[-1.5^n = -7.59325\] divide both by -1 \[1.5^n = 7.59325\] take the log of both sides... doesn't matter which type of log \[\ln(1.5^n) = \ln(7.59325)\] or \[n \ln(1.5) = \ln(7.59325)\] so now solve for n

OpenStudy (campbell_st):

and you'll find n = 5...